not getting the understanding of these problems and they need to be self explanatory to me.
8,11,14,17,20
1,16,81,256,625
5,15,45,135,405
2,7,12,17,22
1,1/2,1/4,1/8,1/16

Answer by jim_thompson5910(35256)

Hints:

8,11,14,17,20 -- adding 3 each time

1,16,81,256,625 -- 1 = 1^4, 16 = 2^4, 81 = 3^4, ...

5,15,45,135,405 -- multiplying by 3 each time

2,7,12,17,22 -- adding 9 each time

1,1/2,1/4,1/8,1/16 -- dividing by 2 each time